OpenStudy (anonymous):

PLEASE HELP? There is a sales tax of $6 on an item that costs $88 before tax. The sales tax on a second item is $21 . How much does the second item cost before tax?

OpenStudy (anonymous):

the tax is \(\large \frac{ 6 }{ 88 }\) on the first item. and the tax is 21 on the second item. so \[\large \frac{ 6 }{ 88 }=\frac{ 21 }{ x }\]

OpenStudy (anonymous):

How much does the second item cost before tax though

OpenStudy (anonymous):

solve for x and you'll know

OpenStudy (anonymous):

okay I got it thank you